Please help me with this geometry question
Pyramid A is a square pyramid with a base side length of 12 inches and a height of 8 inches. Pyramid B has a volume of 3,456 cubic inches. How many times bigger is the volume of pyramid B than pyramid A? Give your answer as a percentage.

Found 3 solutions by mananth, ikleyn, greenestamps:
Answer by mananth(16946)


Volume of square pyramid A
area of base * height
Volume(A) = 12^2* 8= 1152 in^3
Pyramid B has a volume of 3,456 cubic inches.)
(3456/1152) *100
B is three times bigger or 300%

Answer by ikleyn(52824)

.

Looking into the calculations of the volume of the pyramid in the post by @mananth,

notice that all his calculations are INCORRECT:


            ************************************************************************

                the volume of a pyramid IS NOT a product of the base area by the altitude:

                it is only one third of such a product.

            ************************************************************************




Answer by greenestamps(13203)



V(A) = (1/3)(base)(height) = (1/3)(144)(8)=384

V(B) = 3456

V(B)/V(A) = 3456/384 = 9

The volume of pyramid B is 9 times the volume of pyramid A.

However, the answer to the question as asked in the post is 8, not 9, because the question was how many times BIGGER the volume of pyramid B is than the volume of pyramid A.

"9 times as big as" is NOT the same as "9 times bigger than".

"9 times BIGGER THAN x" is x, plus 9 times x, which is 10x.

So, with pyramid B having a volume 9 times AS BIG AS pyramid A, the volume of pyramid B is 8 times BIGGER THAN the volume of pyramid A.

ANSWER: 8